# Langevin equations and a geometric integration scheme for the overdamped limit of homogeneous rotational Brownian motion

### v2

@@ IV.2 Estimation of the numerical error
The form of Eq. 24 admits an exact representation of the solution of the SDE (10) whereas Eq. 26 describes merely an approximation because
$\exp(\Delta\boldsymbol{\Omega}_{A}\cdot\boldsymbol{J})\exp(\Delta\boldsymbol{\Omega}_{B}\cdot\boldsymbol{J})\neq\exp((\Delta\boldsymbol{\Omega}_{A}+\Delta\boldsymbol{\Omega}_{B})\cdot\boldsymbol{J})$ (29)
for $\Delta\boldsymbol{\Omega}_{A}$ and $\Delta\boldsymbol{\Omega}_{B}$ of subsequent integration steps. In order to estimate the discretisation error, we compare the obtained distribution of $\boldsymbol{u}(t+\Delta t)$ with the propagator of the dynamics of $\boldsymbol{u}(t)$, putting $\boldsymbol{T}_{\text{ext}}(t)=0$ for simplicity. Then, the free propagator is known from the solution of the corresponding Fokker–Planck equation and, assuming that $\boldsymbol{u}(0)$ is sampled in equilibrium, it reads in terms of the correlation functions [47]
@@
where $P_{\ell}(\cdot)$ denotes the Legendre polynomial of degree $\ell$. Since $\boldsymbol{u}(t)$ is a Markov process, it is sufficient to consider the first step from $t=0$ to $t=\Delta t$. On the other hand, the numerical scheme [Eq. 25] yields $\boldsymbol{u}(t)\cdot\boldsymbol{u}(0)=\cos(\vartheta)$ and thus
@@
where we have used that $\boldsymbol{n}$ is orthogonal to $\boldsymbol{u}(0)$ and $|\boldsymbol{u}(0)|=1$ and the average is taken with respect to the distribution of $\Delta\boldsymbol{\Omega}$ given in Eq. 28. Inserting the Gaussian distributions for $\Omega_{1}$ and $\Omega_{2}$ and transforming to the angle $\vartheta=(\Omega_{1}^{2}+\Omega_{2}^{2})^{1/2}$, one obtains
@@
This integral can be expressed in terms of Dawson’s $F$-function [85]. For our purposes, however, it suffices to expand the integrand in $\vartheta=0$ since the distribution is sharply peaked for small $D_{\text{R}}t\ll 1$:
@@
×$\qquad\times\frac{1}{4D_{\text{R}}t}\,\text{e}^{-\vartheta^{2}/4D_{\text{R}}t}\,\mathrm{d}(\vartheta^{2})$
@@
where $\alpha_{2}(\ell)=\ell(\ell+1)[3\ell(\ell+1)-2]/8$. From the comparison with Eq. 30 and writing $\Delta t$ again instead of $t$, one finds the relative numerical error of the correlation functions $C_{\ell}(t)$ after one integration step of length $\Delta t$:
@@
asymptotically as $D_{\text{R}}\Delta t\to 0$. In order to integrate $\boldsymbol{u}(0)$ up to time $t=N\Delta t$ one applies the one-step propagator [Eq. 24] successively $N$ times. Correspondingly, the relative discretisation error [Eq. 34] adds up $N$ times so that the overall error obeys
@@ IV.3 A weakly exact integration scheme
A weakly exact integration scheme (i.e., one that yields the correct statistics of the solution) is obtained upon replacing Eq. 28 by the true statistics of $\Delta\boldsymbol{\Omega}$, which tends to a Gaussian only for $D_{\text{R}}\Delta t\ll 1$. In particular, a series representation of the distribution of the rotation angle $\vartheta=|\Delta\boldsymbol{\Omega}|$, restricted to $0\leqslant\vartheta\leqslant\pi$, follows from its angular moments [Eq. 30]:
@@
With this, $C_{\ell}(t)=\int_{0}^{\pi}P_{\ell}\boldsymbol{(}\cos(\vartheta)\boldsymbol{)}\,p(\vartheta,t)\,\mathrm{d}\vartheta$ and reweighting the angle $\vartheta$ in step (iii) of the algorithm in Table 1 such that it samples $p(\vartheta)$ yields numerical correlation functions $\widehat{C}_{\ell}(t)$ [Eq. 31] which are equal to the exact solution $C_{\ell}(t)$ for all $t$. A truncation of the series for $p(\vartheta,t)$ by keeping only terms with $\ell\leqslant\ell_{\text{max}}$ yields $\widehat{C}_{\ell}(t)=C_{\ell}(t)$ for $\ell\leqslant\ell_{\text{max}}$ and $\widehat{C}_{\ell}(t)=0$ otherwise. We note that $p(\vartheta,t)$ is similar in shape, but different from the van-Mises-Fisher distribution, $p_{\text{vMF}}(\vartheta,t)\propto\exp(\cos(\vartheta)/2D_{\text{R}}t)\sin(\vartheta)$, which decays too slowly as $\vartheta\to\pi$ and would yield a larger integration error than given in Eq. 34.
@@
IV.4 Comparison to the ad hoc scheme
Figure 1: Numerically obtained correlation functions $\widehat{C}_{\ell}(t)$ for $\ell=1$ (circles) and $\ell=2$ (diamonds) using the geometric integrator (algorithm in Table 1, closed symbols) and the ad hoc scheme [Eqs. (38), open symbols], which is based on the Euler–Maruyama discretisation of the Langevin equation (7) with an a posteriori normalisation of the orientation vector. Colours distinguish different integration time steps $\Delta t$, numerical data are available only for correlation times $t$ that are a multiple of $\Delta t$, and coloured lines serve as guides to the eye. The solid and dashed black lines show the exact solutions $C_{\ell}(t)$ for $\ell=1$ and $\ell=2$, respectively [Eq. 30].
For completeness, we also give a brief analysis of the numerical scheme suggested in [66], which combines an Euler–Maruyama step for the Stratonovich SDE (7) with a post-normalisation step of the unit vector; the second step corresponds to an orthogonal projection onto the constraining manifold, $|\boldsymbol{u}(t+\Delta t)|=1$. This scheme corresponds to the geometric algorithm in Table 1 with step (iv) replaced by:
@@ IV.4 Comparison to the ad hoc scheme
as before, $\Delta\boldsymbol{\Omega}(t)$ is a Gaussian random vector satisfying Eq. 28. The first equation may also be obtained from Eq. 25 by expanding in $\vartheta=0$ and keeping only terms $O(\vartheta)$.
A calculation similar to the one in Section III.2 shows that the post-normalisation in Eq. 38b is equivalent (in distribution) to adding the Itō drift term to Eq. 38a up to linear order in $\Delta t\to 0$:
@@
We conclude that the ad hoc scheme has the same convergence properties as the Euler–Maruyama scheme for the Itō–Langevin equation [Eq. 9]; in particular, it also exhibits weak convergence of order 1. However, the magnitude of the error is significantly larger compared to the geometric scheme (Table 1). This is demonstrated by numerical results for the correlation functions $\widehat{C}_{1}(t)$ and $\widehat{C}_{2}(t)$ obtained from both schemes (Fig. 1). For an integration time step of $\Delta t=0.1D_{\text{R}}^{-1}$, the results from the ad hoc scheme deviate already considerably from the exact solutions, whereas merely insignificant deviations are visible for $\Delta t=0.3D_{\text{R}}^{-1}$ using the geometric scheme.
Figure 2: Probability distribution $p(\vartheta,\Delta t)$ of the angle $\vartheta$ between the initial and final orientations after an integration time step of length $\Delta t=0.3D_{\text{R}}^{-1}$ for the geometric scheme with Gaussian rotational vectors [Eq. 40, green line] and for the ad hoc scheme, combining an Euler–Maruyama step with an orthogonal projection [Eq. 43, red line]. The grey line shows the exact propagator of rotational Brownian motion [Eq. 37].
For the geometric scheme with Gaussian rotation vectors, the wrapped rotation angle $0\leqslant\vartheta\leqslant\pi$ is distributed with the probability density [see Eq. 32]:
$p_{\text{Gauss}}(\vartheta,\Delta t)$ $=\sum_{k\geqslant 0}\frac{\vartheta+2\pi k}{2D_{\text{R}}\Delta t}\,\text{e}^{-(\vartheta+2\pi k)^{2}/4D_{\text{R}}\Delta t}\,.$ (40)
For most practical purposes, it is sufficient to restrict the sum to the $k=0$ term.
In the case of the ad hoc scheme, the final rotation angle is given via [see Eq. 39]:
@@
$\bar{p}_{\text{ad hoc}}(z,\Delta t)$ $=\int_{0}^{\infty}\delta\left(z-\frac{1}{\sqrt{1+\vartheta^{2}}}\right)\,\frac{\text{e}^{-\vartheta^{2}/4D_{\text{R}}\Delta t}}{2D_{\text{R}}\Delta t}\,\vartheta\mathrm{d}\vartheta$
@@
Zero or negative values, $z\leqslant 0$, cannot be attained in this scheme since $\boldsymbol{u}^{\prime}$ is in the tangent plane of $\boldsymbol{u}$ and thus its projection cannot reach the lower half-sphere. It follows that the probability density of the angle $20\leqslant\vartheta<\pi/$ 2 reads:
$p_{\text{ad hoc}}(\vartheta,\Delta t)=\frac{\sin(\vartheta)}{2D_{\text{R}}\Delta t\cos(\vartheta)^{3}}\,\exp\left(-\frac{\cos(\vartheta)^{-2}-1}{4D_{\text{R}}\Delta t}\right)\,.$ (43)
@@
$p_{\text{ad hoc}}(\vartheta\to 0,\Delta t)=p_{\text{Gauss}}(\vartheta,\Delta t)\,\bigl{[}1+O(\vartheta^{2}/D_{\text{R}}\Delta t)\bigr{]}\,.$ (44)
The obtained distributions $p(\vartheta,\Delta t)$ are the propagators of the various integration schemes investigated. Their graphical comparison in Fig. 2 for $\Delta t=0.3D_{\text{R}}^{-1}$ shows that the geometric scheme with Gaussian rotation vectors (Table 1) closely follows the exact distribution [Eq. 37], despite the comparably large time step. In contrast, the propagator of the ad hoc scheme deviates significantly, which underpins the observations made before for the correlation functions (Fig. 1).
@@ V Extension to axisymmetric particles
where $\mathcal{P}_{\boldsymbol{u}}(t)=\boldsymbol{u}(t)\otimes\boldsymbol{u}(t)$ is the orthogonal projector on the instantaneous direction $\boldsymbol{u}(t)$ and $\mathcal{Q}_{\boldsymbol{u}}(t)=\mathds{1}-\mathcal{P}_{\boldsymbol{u}}(t)$ its complement. For Brownian particles, the argument of time scale separation leading to Eq. 4 applies analogously since the matrix norms of $\boldsymbol{\zeta}_{R}(t)$ and $\boldsymbol{I}(t)$ remain unchanged under the rotational transformations. Thus, in the overdamped limit, Eq. 5 holds with a time-dependent, stochastic friction tensor $\boldsymbol{\zeta}_{R}(t)$. Multiplying the equation with the reciprocal tensor,
@@
$\boldsymbol{\omega}(t)=\boldsymbol{\zeta}_{R}(t)^{-1}\boldsymbol{T}_{\text{ext}}(t)+\boldsymbol{\zeta}_{R}(t)^{-1}\boldsymbol{\xi}(t)\,.$ (48)
For application to the orientation vector, we note that only the component perpendicular to the orientation, $\boldsymbol{\omega}_{\perp}(t)=\mathcal{Q}_{\boldsymbol{u}}(t)\,\boldsymbol{\omega}(t)$, is relevant for the cross product in Eq. 9; it reads
$\boldsymbol{\omega}_{\perp}(t)=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)+(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{\xi}(t)\,,$ (49)
which contains only the friction coefficient $\zeta_{R}^{\perp}$. Also, the evolution of $\boldsymbol{\omega}_{\perp}(t)$ does not depend on the remaining component $\boldsymbol{\omega}_{\parallel}(t)=\mathcal{P}_{\boldsymbol{u}}(t)\,\boldsymbol{\omega}(t)$, parallel to the particle axis. This allows one to replace $\boldsymbol{\omega}(t)$ in Eq. 9 by a process $\tilde{\boldsymbol{\omega}}(t)$, which satisfies Eq. 5 for isotropic friction, i.e., to set $\zeta_{R}^{\parallel}=\zeta_{R}^{\perp}=\zeta_{R}$. As an important consequence, the above discussion and analysis of the overdamped dynamics of $\boldsymbol{u}(t)$ for spherical particles carries over unchanged to axisymmetric particles. In particular, the orientation vector performs a homogeneous Brownian motion on the sphere (for constant $\boldsymbol{T}_{\text{ext}}$) and the described numerical integration scheme can be employed without modifications upon choosing $\zeta_{R}=\zeta_{R}^{\perp}$.
The last point may also be seen from a different angle: since the white noise $\boldsymbol{\xi}(t^{\prime})$ and the orientation $\boldsymbol{u}(t)$ are independent for $t^{\prime}\geqslant t$, the term $\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{\xi}(t)$ in Eq. 49 is equivalent (in probability) to an effectively two-dimensional, Gaussian white noise process that, for each time $t$, is rotated into the plane perpendicular to $\boldsymbol{u}(t)$. This construction is already reflected in the algorithm when choosing the rotation angle in Eq. 27.
For completeness, we specify also the mean and covariance of $\boldsymbol{\omega}_{\perp}(t)$ for $t^{\prime}\geqslant t$, conditioned on the value of $\boldsymbol{u}(t)$ [cf. Eq. 6]:
$\langle\boldsymbol{\omega}_{\perp}(t)|\boldsymbol{u}(t)\rangle$ $=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)\,,$ (50a)

### v3

@@ IV.2 Estimation of the numerical error
The form of Eq.˜24 admits an exact representation of the solution of the SDE (10) in the sense that any given pair of $\boldsymbol{u}(t)$ and $\boldsymbol{u}(t+\Delta t)$ defines a rotation vector $\Delta\boldsymbol{\Omega}(t)$. However,
$\Delta\boldsymbol{\Omega}(t)\neq\int_{t}^{t+\Delta t}\boldsymbol{\omega}(s)\mathrm{d}s$ (29)
in general and, thus, Eq.˜26 describes merely an approximation for finite $\Delta t$; the reason behind being the non-commutativity of finite rotations:
$\text{e}^{\Delta\boldsymbol{\Omega}_{A}\cdot\boldsymbol{J}}\text{e}^{\Delta\boldsymbol{\Omega}_{B}\cdot\boldsymbol{J}}\neq\text{e}^{(\Delta\boldsymbol{\Omega}_{A}+\Delta\boldsymbol{\Omega}_{B})\cdot\boldsymbol{J}}$ (30)
for $\Delta\boldsymbol{\Omega}_{A}$ and $\Delta\boldsymbol{\Omega}_{B}$ of subsequent integration intervals.
In order to estimate the discretization error, we compare the numerically obtained distribution of $\boldsymbol{u}(t+\Delta t)$ with the propagator of the dynamics of $\boldsymbol{u}(t)$, putting $\boldsymbol{T}_{\text{ext}}(t)=0$ for simplicity. The free propagator is known exactly from the solution of the corresponding Fokker–Planck equation [cf. Section˜VI] and, assuming that $\boldsymbol{u}(0)$ is sampled in equilibrium, it reads in terms of the correlation functions [47]
@@
where $P_{\ell}(\cdot)$ denotes the Legendre polynomial of degree $\ell$. Since $\boldsymbol{u}(t)$ is a Markov process, it is sufficient to consider the first step from $t=0$ to $t=\Delta t$. The numerical scheme [Eq.˜25] yields $\boldsymbol{u}(t)\cdot\boldsymbol{u}(0)=\cos(\vartheta)$ and thus for the “numerical” propagator:
@@
where we have used that $\boldsymbol{n}\perp\boldsymbol{u}(0)$ by construction and $|\boldsymbol{u}(0)|=1$; the average is taken with respect to the distribution of $\Delta\boldsymbol{\Omega}$ given in Eq.˜28. Inserting the Gaussian distributions for $\Omega_{1}$ and $\Omega_{2}$ and transforming to the (unwrapped) angle $\vartheta=(\Omega_{1}^{2}+\Omega_{2}^{2})^{1/2}$, one obtains an explicit expression for the propagator corresponding to Eq.˜25:
@@
This integral can be written in terms of Dawson’s $F$-function [92]. For our purposes, however, it suffices to expand the integrand in $\vartheta=0$ since the distribution is sharply peaked for small $D_{\text{R}}t\ll 1$:
@@
$\qquad\times\frac{1}{4D_{\text{R}}t}\,\text{e}^{-\vartheta^{2}/4D_{\text{R}}t}\,\mathrm{d}(\vartheta^{2})$
@@
where $\alpha_{2}(\ell)=\ell(\ell+1)[3\ell(\ell+1)-2]/8$. From the comparison with Eq.˜31 and writing $\Delta t$ again instead of $t$, one finds the relative numerical error of the correlation functions $C_{\ell}(t)$ after one integration step of length $\Delta t$:
@@
asymptotically as $D_{\text{R}}\Delta t\to 0$. In order to integrate $\boldsymbol{u}(0)$ up to time $t=N\Delta t$ one performs the propagation step in Eq.˜24 successively $N$ times. Correspondingly, the relative discretization error [Eq.˜35] adds up $N$ times so that the overall error obeys
@@ IV.3 A weakly exact integration scheme
A weakly exact integration scheme (i.e., one that yields the correct statistics of the solution [51, 63]) is obtained upon replacing Eq.˜28 by the true statistics of $\Delta\boldsymbol{\Omega}$, which tends to a Gaussian only for $D_{\text{R}}\Delta t\ll 1$. In particular, a series representation of the distribution of the rotation angle $\vartheta=|\Delta\boldsymbol{\Omega}|$, here restricted to $0\leqslant\vartheta\leqslant\pi$, follows from its angular moments [Eq.˜31]:
@@
With this, $C_{\ell}(t)=\int_{0}^{\pi}P_{\ell}\boldsymbol{(}\cos(\vartheta)\boldsymbol{)}\,p(\vartheta,t)\,\mathrm{d}\vartheta$ and reweighting the angle $\vartheta$ in step (iii) of the algorithm in Table˜1 such that it samples $p(\vartheta)$ yields numerical correlation functions $\widehat{C}_{\ell}(t)$ [Eq.˜32] which are equal to the exact solution $C_{\ell}(t)$ for all $t$. A truncation of the series for $p(\vartheta,t)$ by keeping only terms with $\ell\leqslant\ell_{\text{max}}$ yields $\widehat{C}_{\ell}(t)=C_{\ell}(t)$ for $\ell\leqslant\ell_{\text{max}}$ and $\widehat{C}_{\ell}(t)=0$ otherwise. We note that $p(\vartheta,t)$ is similar in shape, but different from the van-Mises-Fisher distribution, $p_{\text{vMF}}(\vartheta,t)\propto\exp(\cos(\vartheta)/2D_{\text{R}}t)\sin(\vartheta)$, which decays too slowly as $\vartheta\to\pi$ and would yield a larger integration error than given in Eq.˜35.
@@
IV.4 Comparison to Briels’ projection scheme
Figure 1: Numerically obtained correlation functions $\widehat{C}_{\ell}(t)$ for $\ell=1$ (circles) and $\ell=2$ (diamonds) using the geometric integrator (algorithm in Table˜1, closed symbols) and the projection scheme [Eqs. (39), open symbols], which combines Euler–Maruyama discretization of the Langevin equation (7) with a posteriori normalization of the orientation vector. Colors distinguish different integration time steps $\Delta t$, numerical data are available only for correlation times $t$ that are a multiple of $\Delta t$, and colored lines serve as guides to the eye. The solid and dashed black lines show the exact solutions $C_{\ell}(t)$ for $\ell=1$ and $\ell=2$, respectively [Eq.˜31].
For comparison, we also give a brief analysis of the numerical scheme suggested by Briels et al. [67], which combines an Euler–Maruyama step for the Stratonovich SDE (7) with a post-normalization step of the unit vector; the second step corresponds to an orthogonal projection onto the constraining manifold, $|\boldsymbol{u}(t+\Delta t)|=1$. This scheme corresponds to the geometric algorithm in Table˜1 with step (iv) replaced by:
@@ IV.4 Comparison to Briels’ projection scheme
as before, $\Delta\boldsymbol{\Omega}(t)$ is a Gaussian random vector satisfying Eq.˜28. The first equation may also be obtained from Eq.˜25 by expanding in $\vartheta=0$ and keeping only terms $O(\vartheta)$.
A calculation similar to the one in Section˜III.2 shows that the post-normalization in Eq.˜39b is equivalent (in distribution) to adding the Itō drift term to Eq.˜39a up to linear order in $\Delta t\to 0$:
@@
We conclude that the projection scheme has the same convergence properties as the Euler–Maruyama scheme for the Itō–Langevin equation [Eq.˜9]; in particular, it also exhibits weak convergence of order 1. However, the magnitude of the error is significantly larger compared to the geometric scheme (Table˜1). This is demonstrated by numerical results for the correlation functions $\widehat{C}_{1}(t)$ and $\widehat{C}_{2}(t)$ obtained from both schemes (Fig.˜1). For an integration time step of $\Delta t=0.1D_{\text{R}}^{-1}$, the results from the projection scheme deviate already considerably from the exact solutions, whereas merely insignificant deviations are visible for $\Delta t=0.3D_{\text{R}}^{-1}$ using the geometric scheme.
Figure 2: Probability distribution $p(\vartheta,\Delta t)$ of the angle $\vartheta$ between the initial and final orientations after an integration time step of length $\Delta t=0.3D_{\text{R}}^{-1}$ for the geometric scheme with Gaussian rotational vectors [Eq.˜41, green line] and for Briels’ projection scheme, combining an Euler–Maruyama step with an orthogonal projection [Eq.˜44, red line]. The gray line shows the exact propagator of rotational Brownian motion [Eq.˜38].
For the geometric scheme with Gaussian rotation vectors, the wrapped rotation angle $0\leqslant\vartheta\leqslant\pi$ is distributed with probability density [see Eq.˜33]
$p_{\text{Gauss}}(\vartheta,\Delta t)$ $=\int_{0}^{\infty}\frac{\vartheta^{\prime}}{2D_{\text{R}}\Delta t}\,\text{e}^{-(\vartheta^{\prime})^{2}/4D_{\text{R}}\Delta t}$
$\qquad\times\delta\boldsymbol{(}\cos(\vartheta)-\cos(\vartheta^{\prime})\boldsymbol{)}\,|\sin(\vartheta^{\prime})|\mathrm{d}\vartheta^{\prime}\,.$
$=\sum_{k=0}^{\infty}\frac{\vartheta+2\pi k}{2D_{\text{R}}\Delta t}\,\text{e}^{-(\vartheta+2\pi k)^{2}/4D_{\text{R}}\Delta t}$
$\qquad+\sum_{k=1}^{\infty}\frac{2\pi k-\vartheta}{2D_{\text{R}}\Delta t}\,\text{e}^{-(2\pi k-\vartheta)^{2}/4D_{\text{R}}\Delta t}\,.$ (41)
For most practical purposes, it is sufficient to restrict the sum to the $k=0$ term; the next-to-leading order term ($k=1$) is suppressed by a factor of $\exp(-\pi^{2}/4D_{\text{R}}\Delta t)$.
In the case of the projection scheme, the final rotation angle is given via [see Eq.˜40]:
@@
$\bar{p}_{\text{proj}}(z,\Delta t)$ $=\int_{0}^{\infty}\delta\left(z-\frac{1}{\sqrt{1+\vartheta^{2}}}\right)\,\frac{\text{e}^{-\vartheta^{2}/4D_{\text{R}}\Delta t}}{2D_{\text{R}}\Delta t}\,\vartheta\mathrm{d}\vartheta$
@@
Zero or negative values, $z\leqslant 0$, cannot be attained in this scheme since $\boldsymbol{u}^{\prime}$ is in the tangent plane of the constraint at point $\boldsymbol{u}$ and thus its projection cannot reach the lower half-sphere. It follows that the probability density of the angle $0\leqslant\vartheta<\pi/2$ reads:
$p_{\text{proj}}(\vartheta,\Delta t)=\frac{\sin(\vartheta)}{2D_{\text{R}}\Delta t\cos(\vartheta)^{3}}\,\exp\left(-\frac{\cos(\vartheta)^{-2}-1}{4D_{\text{R}}\Delta t}\right)\,.$ (44)
@@
$p_{\text{proj}}(\vartheta\to 0,\Delta t)=p_{\text{Gauss}}(\vartheta,\Delta t)\,\bigl{[}1+O(\vartheta^{2}/D_{\text{R}}\Delta t)\bigr{]}\,.$ (45)
The obtained distributions $p(\vartheta,\Delta t)$ are the propagators of the various integration schemes investigated. Their graphical comparison in Fig.˜2 for $\Delta t=0.3D_{\text{R}}^{-1}$ shows that the geometric scheme with Gaussian rotation vectors (Table˜1) closely follows the exact distribution [Eq.˜38], despite the comparably large time step. In contrast, the propagator of the projection scheme deviates significantly, which underpins the observations made before for the correlation functions (Fig.˜1).
@@ V Extension to axisymmetric particles
where $\mathcal{P}_{\boldsymbol{u}}=\boldsymbol{u}\otimes\boldsymbol{u}$ is the orthogonal projector on the given direction $\boldsymbol{u}$ and $\mathcal{Q}_{\boldsymbol{u}}=\mathds{1}-\mathcal{P}_{\boldsymbol{u}}$ its complement. For Brownian particles, the argument of time scale separation leading to Eq.˜4 applies analogously since the matrix norms of $\boldsymbol{\zeta}_{R}(t)$ and $\boldsymbol{I}(t)$ remain unchanged under the rotational transformations. Thus, in the overdamped limit, Eq.˜5 holds with a time-dependent, stochastic friction tensor $\boldsymbol{\zeta}_{R}(t)$. Multiplying the equation with the reciprocal tensor,
@@
$\boldsymbol{\omega}(t)=\boldsymbol{\zeta}_{R}(t)^{-1}\boldsymbol{T}_{\text{ext}}(\boldsymbol{u}(t),t)+\boldsymbol{\zeta}_{R}(t)^{-1}\boldsymbol{\xi}(t)\,.$ (49)
Here, we have also included a possible dependence of the external torque on the instantaneous orientation $\boldsymbol{u}(t)$, which is commonly encountered in physical systems, e.g., with dipolar interactions.
For application to the orientation vector, one observes that only the component perpendicular to the orientation, $\boldsymbol{\omega}_{\perp}(t)=\mathcal{Q}_{\boldsymbol{u}(t)}\,\boldsymbol{\omega}(t)$, is relevant for the cross product in Eq.˜9; it reads
$\boldsymbol{\omega}_{\perp}(t)=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}(t)}\boldsymbol{T}_{\text{ext}}(\boldsymbol{u}(t),t)\\ +(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}(t)}\boldsymbol{\xi}(t)\,$, (50)
which contains only the friction coefficient $\zeta_{R}^{\perp}$. Also, the evolution of $\boldsymbol{\omega}_{\perp}(t)$ does not depend on the remaining component $\boldsymbol{\omega}_{\parallel}(t)=\mathcal{P}_{\boldsymbol{u}(t)}\,\boldsymbol{\omega}(t)$, parallel to the particle axis. This allows one to replace $\boldsymbol{\omega}(t)$ in Eq.˜9 by a process $\tilde{\boldsymbol{\omega}}(t)$, which satisfies Eq.˜5 for isotropic friction, i.e., to set $\zeta_{R}^{\parallel}=\zeta_{R}^{\perp}=\zeta_{R}$. As an important consequence, the above discussion and analysis of the overdamped dynamics of $\boldsymbol{u}(t)$ for spherical particles carries over unchanged to axisymmetric particles. In particular, the orientation vector performs a homogeneous Brownian motion on the sphere (for time-independent $\boldsymbol{T}_{\text{ext}}$) and the described numerical integration scheme can be employed without modifications upon choosing $\zeta_{R}=\zeta_{R}^{\perp}$.
The last point may also be seen from a different angle: since the white noise $\boldsymbol{\xi}(t^{\prime})$ and the orientation $\boldsymbol{u}(t)$ are independent for $t^{\prime}\geqslant t$, the term $\mathcal{Q}_{\boldsymbol{u}(t)}\boldsymbol{\xi}(t)$ in Eq.˜50 is equivalent (in probability) to an effectively two-dimensional, Gaussian white noise process that, for each time $t$, is rotated into the plane perpendicular to $\boldsymbol{u}(t)$. This construction is already reflected in the algorithm when choosing the rotation angle in Eq.˜27. We note that the algorithm is also applicable if $\boldsymbol{T}_{\text{ext}}(\boldsymbol{u},t)$ depends explicitly on time; however, the error estimates in Section˜IV are no longer valid in this case.
For completeness, we specify the mean and covariance of $\boldsymbol{\omega}_{\perp}(t)$ for $t^{\prime}\geqslant t$, conditioned on the value of $\boldsymbol{u}(t)$ [cf. Eq.˜6]:
$\langle\boldsymbol{\omega}_{\perp}(t)|\boldsymbol{u}(t)\rangle$ $=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}(t)}\boldsymbol{T}_{\text{ext}}(\boldsymbol{u}(t),t)\,$, (51a)